# Continual Adversarial Defense

### v4

@@ Introduction
Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and OnePixel($l_{0}$) to validate the efficacy of each component. As shown in Table 3, CAD without $\mathcal{L}_{0}$ and $\mathcal{L}_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage. The absence of model ensemble leads to the inability to uphold performance on clean images.
Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Figure 3 to visualize the defense capability of CAD after adaptation. These features are encoded by the defense model after the last stage of adapting to OnePixel attack. From (a) to (d) in Figure 3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types. After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
Multiple Attacks in One Stage
Stage 0 1 2 3
$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
Table 4: Acc(%) against multiple attacks in one stage.
In practice, it is common for multiple attacks to occur simultaneously within a single stage. To assess CAD’s adaptability in such scenarios, on CIFAR-10, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks per stage. As presented in Table 4, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
Number of Shots for Adaptation
Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel
$K=15$ 94.5 90.0 88.4 85.5
$K=10$ 94.5 89.9 86.6 85.3
$K=5$ 94.5 89.4 86.1 84.6
$K=1$ 94.5 87.9 84.6 83.4
Table 5: AIAcc(%) with different budget scale.
@@
Memory Allowance for Adaptation
As outlined in Principle 3, we need to employ memory efficiently for continual defense. In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each
Cache Size Model Size
Stage 1 2 3 1 2 3
@@
Table 6: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD.
@@
stage on CIFAR-10. As depicted in Table 6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
Conclusion
@@
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect few-shot defense budget from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles that include continual adaptation without catastrophic forgetting, few-shot budget adaptation, memory-efficient adaptation, and high accuracy for both clean and adversarial data, and CAD demonstrates strong performance when operating under these principles. Through extensive experiments, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
Appendix
Algorithm 1 Continual Adversarial Defense
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot budget $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.
Output: Defense model ensemble.
1:  Generate initial adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}$ using $A_{0}$.
2:  Train weight estimator $f_{p}$ on $\mathcal{D}_{\mathrm{train}}$ using self-perturbation.
3:  Rewrite the ground-truth of $\mathcal{A}_{\mathrm{train}}^{i}$ using Eq.1.
4:  Optimize defense model $f_{0}$ on $\mathcal{A}_{\mathrm{train}}^{0}$ using Eq.4.
5:  Generate prototypes of $A_{0}$ using Eq.7 and add them to prototype set $P_{e}$.
6:  for $i$ in $1,\cdots,T$ do
7:     Fine-tune the part of the classifier preserved for the new attack using Eq.8.
8:     Generate prototypes of $A_{i}$ using Eq.7 and add them to prototype set $P_{e}$.
9:     Refine the classifier using cached prototypes $P_{e}$.
10:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq.9.
11:  end for
Appendix A More Details about CAD
The overall algorithm of the CAD framework is presented in Algorithm 1.
Model Ensemble Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation (Wang et al. 2023a) to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.
Algorithm 2 Self-perturbation
Input: A empty perturbation matrix $\boldsymbol{\eta}^{p}\in\mathbb{R}^{H\times W\times 3}$ with the same shape of real image $\mathbf{x}^{r}$.
Parameter: Max perturbation magnitude $\epsilon$, pattern mode.
Output: Self-perturbed image $\mathbf{x}^{p}$.
1:  A random direction matrix $R=\{\vec{r}_{ij}\}\in\mathbb{R}^{H\times W\times 3}$.
2:  for $\vec{\eta}_{ij}$ in $\boldsymbol{\eta}^{p}$ do
3:     Select random noise value $\alpha$.
4:     if pattern mode is ‘point-wise’ then
5:        $\vec{\eta}_{ij}:=\alpha\cdot\vec{r}_{ij}$.
6:     else if pattern mode is ‘block-wise’ then
7:        Select a random neighborhood $A_{ij}$ of $\vec{\eta}_{ij}$.
8:        for $\vec{\eta}_{ijk}$ in $A_{ij}$ do
9:           $\vec{\eta}_{ijk}:=\alpha\cdot\vec{r}_{ij}$.
10:        end for
11:     end if
12:  end for
13:  Clip perturbation $\boldsymbol{\eta}^{p}$ using Equation 10.
14:  Generate self-perturbed image $\mathbf{x}^{p}$ using Equation 11.
15:  return $\mathbf{x}^{p}$
As presented in Algorithm 2, we perturb each point in the point-wise pattern and each block in the block-wise pattern in a stochastic direction, where blocks are random neighborhoods of a set of scattered points. The generated perturbation image $\boldsymbol{\eta}^{p}$ is constrained in $l^{\infty}$ norm, and clipped according to $\epsilon$,
$\boldsymbol{\eta}^{p}=\mathrm{Clip}_{[-\epsilon,\epsilon]}(\boldsymbol{\eta}^{p}).$ (10)
A self-perturbed image is calculated as
$\mathbf{x}^{p}=\mathbf{x}^{r}+\boldsymbol{\eta}^{p}.$ (11)
Trained with self-perturbation, the weight estimator $f_{p}$ is able to distinguish between the clean and adversarial images. The output logit $\mathbf{w}$ of $f_{p}$ represents the probability of an image to be clean or adversarial. Therefore, we use $\mathbf{w}$ to ensemble the target model and the defense model:
In this way, both the clean images and the adversarial images are correctly classified.
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
TRADES (Zhang et al. 2019) 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
JEM (Grathwohl et al. 2020) 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
PAT (Laidlaw, Singla, and Feizi 2020) 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
GAIRAT (Zhang et al. 2020) 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
FastAdv (Jiang et al. 2023) 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
RPF (Dong and Xu 2023) 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
DMAT (Wang et al. 2023b) 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
EBM (Hill, Mitchell, and Zhu 2020) 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
ADP (Yoon, Hwang, and Lee 2021) 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
DiffPure (Nie et al. 2022) 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.6 94.1 94.6 94.3 95.1 95.1 95.7 96.0
Table 7: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense budget.
Appendix B Premium-CAD
We propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In each stage, we train an attack-specific defense model using abundant feedback and ensemble the target model and all the defense models by a stage-evolved weight estimator.
Attack-Specific Defense Model
In stage 00, the initial defense model is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. In stage $i$, another defense model is trained using abundant feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)\}_{N\times K}$ where $N$ is the number of classes in dataset $\mathcal{D}_{\mathrm{train}}$ and $K$ is the amount of feedback per class. The scale of feedback is set to $K=10$ in CAD and $K=1000$ in premium-CAD.
In this way, we have $i+1$ defense models in stage $i$. The defense model is designed to tackle adversarial examples from a specific attack, complementing each other.
Algorithm 3 Premium Continual Adversarial Defense
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, abundant feedback $\mathcal{A}_{\mathrm{train}}^{i}$ generated by attack $A_{i}$ for stage $i=1,\cdots,T$.
Output: Defense model ensemble.
1:  for $i$ in $1,\cdots,T$ do
2:     Train an attack-specific defense model $f_{i}$.
3:     Train a stage-evolved weight estimator $f_{p}^{i}$.
4:     Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Equation 14.
5:  end for
Stage 0 1 2 3 4 5 6 7 8
Attack VMIM VNIM NIM DIM MIM RFGSM BIM SNIM PGD
CIFAR 94.8 94.6 93.4 93.5 94.2 93.4 92.5 92.2 93.3
ImageNet 82.5 82.1 82.0 82.0 92.4 82.2 81.9 82.3 81.7
Table 8: Classification Acc (%) against changed attack sequence.
Stage-Evolved Weight Estimator
To ensemble the target model $f_{t}$ and defense models $[f_{0},f_{1},\cdots,f_{i}]$, we use a joint dataset $\mathcal{D}_{J}^{i}=\bigcup_{i=0,1,\cdots,i}\mathcal{A}_{\mathrm{train}}^{i}\cup\mathcal{D}_{\mathrm{train}}$ to train the weight estimator $f_{p}^{i}$ in stage $i$. Feed an image to the weight estimator $f_{p}^{i}$, it output a weight vector $\mathbf{w}^{i}\in\mathbb{R}^{i+2}$ to ensemble models:
$\mathbf{w}^{i}=f_{i}(\mathbf{x})$ (13)
The final output for premium-CAD is:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}^{i}\cdot[f_{t}(\mathbf{x}),f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\cdots,f_{i}(\mathbf{x})]^{\top}$ (14)
The overall algorithm of premium-CAD framework is presented in Algorithm 3. The loss function for training attack-specific defense models and stage-evolved weight estimators is Cross-Entropy Loss.
Appendix C More Implementation Details
Adversarial Attack
In this work, all adversarial attacks are implemented using TorchAttacks (Kim 2020). The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al. 2019). Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively. The number of perturbed pixels of OnePixel (Su, Vargas, and Sakurai 2019) is set to $50$. The other parameters are the default ones in TorchAttacks.
Defense Baselines
The adversarial training baselines are implemented using RobustBench (Croce et al. 2021), while purification baselines are implemented using open-source code and model checkpoints. CL baselines are implemented using PyCIL (Zhou et al. 2023). The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100. CL models are trained in the same settings as in CAD. Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.
Appendix D More Experimental Results
Continual Adversarial Defense on CIFAR-10
We compare our CAD with more adversarial defense baselines on CIFAR-10. As presented in Table 7, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT (Wang et al. 2023b) and ADP (Yoon, Hwang, and Lee 2021), CAD outperforms them by approximately $10\%$ for each attack. As for clean images, CAD achieves $95.8\%$ accuracy while the target model stands at $96.4\%$.
Change the sequential of attacks
For different attack sequences, we reverse the sequential of attacks and conduct an experiment. As shown in Table 8, when the attack sequence changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack order.
Change the intensity of attacks
Stage 0 1 2 3 4 5 6
ImageNet-$\epsilon/255$ $4$ $6$ $8$ $10$ $12$ $14$ $16$
Acc 83.7 83.6 83.3 83.3 83.4 82.2 82.5
CIFAR-$\epsilon/255$ $8$ $10$ $12$ $14$ $16$ $18$ $20$
Acc 94.5 94.4 94.4 93.9 94.0 93.2 93.4
Table 9: Classification Acc (%) against various attack intensity of PGD($l_{\infty}$).
For different attack intensities, we conduct experiments using PGD($l_{\infty}$) with various $\epsilon$. As shown in Table 9, when the attack intensity changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack intensity.
Averaged accuracy when the number of classes increases
We present curves depicting the relationship between averaged accuracy and the number of classes in Figure 4. As the number of classes increases, CAD continues to perform well on both the new attack and previous attacks.
Figure 4: Average Acc with 10 classes and 100 classes per stage for CIFAR10 and ImageNet100.
Appendix E Limitations and Future Work
The increase in attacks denotes the expansion of the defense model classifier, especially when tackling large datasets such as the original ImageNet. Although CAD refines the defense budget before each stage’s classifier expansion, too many classes can lead to an undesirable decrease in classification performance. In the future, we will focus on this issue and refine the classifier to guarantee the consistent performance of the defense model.

### v5

@@ 5.1 Comparisons to Baselines
5.2 Ablation Study
As argued, embedding reservation facilitates smoother adaptation and reduces overfitting, prototype augmentation preserves decision boundaries from previous stages, and model ensembling ensures robust classification for both clean and adversarial images. To validate the contribution of each component, we conduct an ablation study using four attacks: PGD ($l_{\infty}$), CW ($l_{2}$), EAD ($l_{1}$), and OnePixel ($l_{0}$). As shown in Table 3, removing $\mathcal{L}_{0}$ and $\mathcal{L}_{\mathrm{proto}}$ results in performance drops of $13.5\%$ and $6.8\%$, respectively, in the final stage. Additionally, removing model ensembling (w/o $f_{p}$) severely degrades performance on clean images. These results confirm the effectiveness of each component.
Figure 4: Catastrophic forgetting verification of CAD and the vanilla model under the attack sequence in Table 2 on CIFAR-10.
5.3 Adaptation Without Forgetting
As stated in Principle 1, a defense model must adapt to new attacks while preventing catastrophic forgetting. To evaluate CAD’s adaptability, we present T-SNE visualizations of features extracted from adversarial test data in Figure 3. These features are encoded by the defense model after the final stage of the attack sequence in Table 2. From (a) to (d) in Figure 3, the feature clusters become increasingly indistinct, indicating that the difficulty of few-shot adaptation grows as new attack types are introduced. Despite this, CAD successfully retains its defense performance against previous attacks, demonstrating its ability to adapt while mitigating catastrophic forgetting.
Furthermore, we compare CAD with a vanilla model as the empirical lower bound. As shown in Figure 4, without CAD, the model’s forgetting of previous attacks worsens over stages,with its error rate on PGD and CW exceeding 50% by stage 8. This highlights CAD’s effectiveness in preventing catastrophic forgetting.
5.4 Number of Shots for Adaptation
$K$ 0:PGD 1:CW 2:EAD 3:OnePixel
$15$ 94.5 90.0 88.4 85.5
$10$ 94.5 89.9 86.6 85.3
$5$ 94.5 89.4 86.1 84.6
$1$ 94.5 87.9 84.6 83.4
Table 4: AIAcc(%) at different budget scale $K$.
@@ 5.4 Number of Shots for Adaptation
Method Cache Size per Stage Model Size per Stage
1 2 3 1 2 3
@@
Table 5: Comparison of the Memory cost (Byte) between CAD and premium-CAD.
5.5 Memory Allowance for Adaptation
@@
As outlined in Principle 3, efficient memory usage is crucial for continual defense. To evaluate CAD’s efficiency, we measure memory footprint and model size at each stage on CIFAR-10. As shown in Table 5, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
5.6 Defense Against White-Box Attack Sequence
To further evaluate the robustness of CAD, we conducted white-box experiments on CIFAR-10. The white-box setting follows the same configuration as AIR [44], where attacks occur in real-time during both adaptation and testing, and there is no restriction on the amount of defense budget used for adaptation. Moreover, both the target model and the defense model are fully visible to the attacker. To adapt to white-box attacks, the feature extractor $\varphi$ is no longer fixed but learnable.
We compare CAD with vanilla adversarial training (AT) and CAD + robust model under a white-box attack sequence. Here, CAD + robust model refers to using a robust model [35] as the backbone in CAD. As shown in Table 6, CAD outperforms vanilla AT in adapting to new attacks and further strengthens the robust model’s defense against evolving threats.
Method 0:PGD 1:SNIM 2:BIM 3:RFGSM
Vanilla AT [22] 36.2 35.8 34.5 34.2
CAD 51.7 55.2 53.6 52.5
$\mathrm{\textbf{CAD}}^{\dagger}$ 51.7 55.4 53.8 53.4
Robust Model [35] 72.1 78.8 71.7 71.6
CAD + Robust Model 74.5 79.6 73.4 73.8
$\mathrm{\textbf{CAD}}^{\dagger}$ + Robust Model 74.5 81.2 78.8 78.9
Table 6: Classification accuracy (%) against white-box attacks.
5.7 Comparison with AIR
We compare CAD’s performance in the white-box setting with the results reported in the AIR paper. As shown in Table 7, after adapting from Attack 1 to Attack 2, CAD better retains its defense against previous attacks compared to AIR. This is attributed to CAD’s prototype augmentation mechanism, albeit at the cost of a slightly larger prototype cache. Additionally, since premium-CAD trains a separate defense model for each attack, it completely avoids forgetting previous attacks after adaptation.
Method FGSM to PGD PGD to FGSM
Attack 1 Attack 2 Attack 1 Attack 2
AIR [44] ${53.4}^{\ddagger}$ ${44.1}^{\ddagger}$ ${43.0}^{\ddagger}$ ${52.3}^{\ddagger}$
CAD 54.2 48.6 47.5 53.9
$\mathrm{\textbf{CAD}}^{\dagger}$ 55.8 51.7 51.7 55.8
[…]
6 Conclusion
@@
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect defense budget from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles that include continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data, and CAD demonstrates strong performance when operating under these principles. Through extensive experiments, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.